OpenStudy (anonymous):

I just verify my solution for a problem below: Obtain the centre and radius of the circle x^2 + y^2 — 6x —12y —19 = 0. Can anybody help me?

OpenStudy (anonymous):

Proceed by completing the square. From there, what did you get and I can verify it :P

OpenStudy (anonymous):

My answer is radius : 8 and cneter is at (3,6)

OpenStudy (anonymous):

Okay, just 2 seconds :P

OpenStudy (jit4won):

center = 6/2 . 12/2= 3,6

OpenStudy (jit4won):

radius=8

OpenStudy (anonymous):

I got: \[(x-3)^2+(y-6)^2=64\]. So yes, radius 8 and center (3,6)

OpenStudy (anonymous):

Thanks for your help.I appreciate it.

OpenStudy (anonymous):

No problem :P

OpenStudy (jit4won):

if you have an equation in the form of x^2 + Y^2 +2gx +2fy+c= 0 use center = -g,-f

OpenStudy (anonymous):

Oh,that's new for me.Thanks jit4won.

OpenStudy (jit4won):

radius = square root of ( g^2+f^2 - c)

OpenStudy (jit4won):

Its a shortcut method ..

OpenStudy (jit4won):

please check it now

OpenStudy (jit4won):

center =( - half of coefficient of x, - half of coefficient of y)

OpenStudy (jit4won):

I can help you if you have skype/gtalk /yahoo/ any other site where i could instruct you